OpenStudy (anonymous):

I'm having trouble solving this problem... -t/4=8??? Am I supposed to multiply it by the reciprocal on both sides??? If so, how do I multiply -t/4 by its reciprocal when there's only 1 number??? I'm so confused!!!

OpenStudy (unklerhaukus):

\[{-t \over 4}=8\]multiply both sides of the equation by 4\[\rightarrow{-t\over 4} \times 4 =8 \times 4\]simplify \[-t=32\] multiply/divide both sides by negative 1 \[-t \times -1=32 \times -1\] simplify \[t=-32\]

OpenStudy (akshay_budhkar):

i need to go to bed @unkle.. my bad sorry :P

OpenStudy (unklerhaukus):

you said it right akshay,

OpenStudy (akshay_budhkar):

lol that was the biggest blunder i could ever do :P . i think i must throw all my math certificates :P

OpenStudy (unklerhaukus):

yeah well i make simple mistakes all the time ,

OpenStudy (akshay_budhkar):

hehe.. that gives me relief :P